A 1.00 g sample of zinc (Zn) dissolves completely in 100.0 mL of hydrochloric acid in a calorimeter (P=const, no heat loss). The solution temperature changes from 22.8 to 28.4 ◦C. Find ∆H◦ rxn, kJ/mol, of the reaction Zn(s) + 2HCl(aq) = ZnCl2(aq) + H2(g).

M(Zn) = 65.38 g/mol, Cs(solution) = 4.18 J/(g· ◦C), d(solution) = 1.00 g/mL. Consider changes of the solution volume negligible.

(A) −153 kJ/mol

(B) −35.80 kJ/mol

(C) −2.341 kJ/mol

(D) +35.80 kJ/mol

(E) +153 kJ/mol
Answer #1

(E) +153 kJ/mole

Zn (s) + 2 HCl (aq) ------------> ZnCl2 (aq) + H2 (g)

Mass of Zn = 1.00 g

Moles of Zn = mass/MW = 1.00 g/65.38 g/mole = 0.015295 mole

Volume = 100.0 mL

Density = 1.00 g/mL

Mass = Volume x Density = 100.0 mL x 1.00 g/mL = 100.0 g

∆ T = 28.4 oC - 22.8 oC = 5.6 oC

We know that,

∆Hrxn = m x Cs x ∆ T

∆Hrxn = 100.0 g x 4.18 J/gC x 5.6 oC

∆Hrxn = 2340.8 J

∆Hrxn in KJ/mole

∆Hrxn = 2340.8 J / 0.015295 mole

∆Hrxn = 153043.478 J/mole

∆Hrxn = 153.04 kJ/mole